# TOTEM: TOkenized Time Series EMbeddings for General Time Series Analysis

### v1

@@ Abstract
The field of general time series analysis has recently begun to explore unified modeling, where a common architectural backbone can be retrained on a specific task for a specific dataset. In this work, we approach unification from a complementary vantage point: unification across tasks and domains. To this end, we explore the impact of discrete, learnt, time series data representations that enable generalist, cross-domain training. Our method, TOTEM, or TOkenized Time Series EMbeddings, proposes a simple tokenizer architecture that embeds time series data from varying domains using a discrete vectorized representation learned in a self-supervised manner. TOTEM works across multiple tasks and domains with minimal to no tuning. We study the efficacy of TOTEM with an extensive evaluation on 17 real world time series datasets across 3 tasks. We evaluate both the specialist (i.e., training a model on each domain) and generalist (i.e., training a single model on many domains) settings, and show that TOTEM matches or outperforms previous best methods on several popular benchmarks. The code can be found at: https://github.com/SaberaTalukder/TOTEM.
representation learning, tokenization, time series
@@ 1 Introduction
Figure 1: TOTEM & Evaluation Regimes. An overview of TOTEM, the training schemas, and inference regimes. (a) The TOTEM VQVAE architecture consists of an 1D strided CNN encoder $\mathcal{E}$, quantizer, latent codebook, and 1D strided transpose CNN decoder $\mathcal{D}$. TOTEM’s VQVAE enables generalist training, i.e., on all datasets jointly, and specialist training, i.e., on one dataset at a time. (b) TOTEM’s discrete, self-supervised codebook can be leveraged for both in domain and zero shot testing.
@@
Time series analysis encompasses a wide range of tasks, datasets, and applications in the real world. When considering training paradigms, time series analysis has historically been conducted via specialist-training, meaning that models are trained on a single time series domain (Zhou et al., 2023; Wu et al., 2022a; Nie et al., 2022; Zhang & Yan, 2022). Generalist-training, where models are simultaneously trained on multiple time series domains, contrasts the specialist paradigm. Both specialist and generalist models can be tested under various regimes. Within in-domain-testing, a model is tested on the same domain(s) it was trained on. In zero-shot-testing, a model is tested on different domains(s) than it was trained on. Some methods have begun to explore the idea of zero-shot forecasting where (1) a forecaster trains on one dataset then predicts on a separate dataset (Zhou et al., 2023), or (2) a forecaster trains on a subset of channels (which we call sensors) from one dataset then zero-shot forecasts on the remaining sensors in the same dataset (Liu et al., 2023). Both of these models would be considered specialists, as they were trained on only one (or a subset of one) dataset.
Further, time series analysis has typically been restricted by task, where methods study only forecasting (Wu et al., 2021; Woo et al., 2022), anomaly detection (Xu et al., 2021; He & Zhao, 2019), or imputation (Luo et al., 2018, 2019; Talukder et al., 2022), among others. Recently, the field has become increasingly unified with respect to model architecture, with methods (Zhou et al., 2023; Wu et al., 2022a) exploring language and vision backbones on various time series tasks. These backbones, like previous methods, utilize specialist training (e.g., training separate imputers on each dataset).
Similarly, the field has also become increasingly unified with respect to data representation, with growing emphasis on learning performant data representations. For instance, Franceschi et al. (2019) utilize an exponentially dilated causal convolutional encoder to discover in-domain embeddings, Tonekaboni et al. (2021) leverage temporal neighborhood coding, Yang & Hong (2022) utilize temporal-spectral fusion, and Yue et al. (2022) employs hierarchical contrasting across the time and batch dimensions.
At a technical level, our approach bears closest affinity to methods that use vector quantized variational autoencoders (VQVAEs) (Van Den Oord et al., 2017; Duan et al., 2023; Rasul et al., 2022b, a). As we discuss further in Section 3, our goal is to develop a streamlined framework for learning a tokenized data representation (using VQVAEs) in a way that permits easy applicability and holistic empirical evaluation on a broad range of time series modeling tasks and data domains (including zero-shot generalization to new test domains) with minimal to no tuning. [^2] As an aside, our approach to studying what is a performant general time series data representation shares a philosophical alignment with the development of large generalist models in natural language processing, which are also based on having a common tokenized representation (Gage, 1994; Radford et al., 2018).
Motivated by the trend of time series analysis unification, we explore the value of a VQVAE-based tokenizer for time series imputation, anomaly detection, and forecasting. Unlike previous methods, we utilize self-supervised, discrete tokens, and extensively explore their utility in varied training and testing regimes. Our contributions are as follows:
  1. 1.
    We present TOTEM, a simple tokenizer architecture for time series analysis that works across domains and tasks with minimal to no tuning.
  2. 2.
    Despite its simplicity, TOTEM matches or outperforms the state-of-the-art on several popular benchmark datasets and tasks.
  3. 3.
    With an extensive evaluation in the generalist setting (training a single model on multiple domains), we show that TOTEM outperforms the leading state-of-the-art model in both in-domain and zero-shot testing regimes.
@@ 2 Related Work
Time series modeling methods utilize many techniques, ranging from statistical methods (Winters, 1960; Holt, 1957; Anderson, 1976; Hyndman & Athanasopoulos, 2018; Taylor & Letham, 2018) to multilayer perceptrons (MLPs) (Zeng et al., 2023; Li et al., 2023; Das et al., 2023; Challu et al., 2023; Chen et al., 2023; Zhang et al., 2022; Oreshkin et al., 2019) to convolutional neural networks (CNNs) (Wu et al., 2022a; Liu et al., 2022a; He & Zhao, 2019; Franceschi et al., 2019; Bai et al., 2018) to recurrent neural networks (RNNs) (Salinas et al., 2020; Shen et al., 2020; Hochreiter & Schmidhuber, 1997) to transformers (Zhou et al., 2023; Liu et al., 2023; Nie et al., 2022; Zhang & Yan, 2022; Woo et al., 2022; Zhou et al., 2022; Liu et al., 2022b; Wu et al., 2022b; Xu et al., 2021; Wu et al., 2021; Liu et al., 2021; Zhou et al., 2021; Kitaev et al., 2020; Li et al., 2019). Many models are hybrid solutions that blend aforementioned approaches.
Most of these methods intake time and then perform various combinations of normalization (Kim et al., 2021), frequency transformations (Wu et al., 2022a; Zhou et al., 2022), and patchification either along the time dimension (Liu et al., 2023; Zhang & Yan, 2022; Nie et al., 2022), or sensor dimension (Li et al., 2019; Zhou et al., 2021; Wu et al., 2021; Liu et al., 2021). [^3] In time series analysis, sensors, channels, and variates are synonymous terms; in this paper we adopt the sensor terminology. Patch lengths range from a single time-step / sensor, also known as point-wise, to the length of the entire time series / all sensors. Time and sensor patch dependencies are then learned, via an attention mechanism, convolution, recurrence, or linear layer, across the temporal dimension, sensor dimension, or both the temporal and sensor dimensions (Zhang & Yan, 2022). For multisensor modeling, one can model all sensors jointly or independently (i.e., forecast each sensor independently) (Nie et al., 2022). These methods learn the underlying data representations end-to-end with the downstream task (e.g., forecasting).
Specialist-training, where models are only trained on a single time series domain, is the most common regime amongst prior work (Zhou et al., 2023; Wu et al., 2022a; Nie et al., 2022; Zhang & Yan, 2022). These specialist models are primarily evaluated via in-domain-testing, where the test set is from the same domain as the train set. Recently, some methods (Zhou et al., 2023; Liu et al., 2023) have begun to explore specialist zero-shot forecasting capabilities.
The time series analysis field is undergoing unification along both the modeling axis (Zhou et al., 2023; Wu et al., 2022a) and data representation axis (Franceschi et al., 2019; Tonekaboni et al., 2021; Yang & Hong, 2022; Barnum et al., 2020; Yue et al., 2022). Unified data representations, both statistical and learnt, have been more extensively studied in language and vision modeling (Gage, 1994; Van Den Oord et al., 2017; Esser et al., 2021; Rombach et al., 2022). The vision modeling field distinguishes between discrete, learnt, tokens (Van Den Oord et al., 2017; Esser et al., 2021; Rombach et al., 2022) and patches (Dosovitskiy et al., 2020). Patches have been studied in time series modeling (Zhou et al., 2023; Nie et al., 2022; Zhang & Yan, 2022). In this work, we propose to use discrete, learnt tokenized representations, which we show lead to strong performance in both specialist and generalist settings, as well as in-domain and zero-shot testing regimes.

### v2

@@ Abstract
This work studies the problem of time series analysis with generalist (or foundation) models, which are models trained across many data domains. Drawing inspiration from the widespread success of large language models, we consider the simple strategy of discretely tokenizing time series data drawn from a myriad of datasets via self-supervision, then using the fixed tokenization to solve a variety of tasks across many data domains. Canonically, time series models are either trained on a single dataset or built in a task-specific manner (e.g., a forecasting-only model), where many use patches of time as inputs to the model. As such, performant generalist, discrete representation time series models explored across many tasks are of value. Our method, TOkenized Time Series EMbeddings (TOTEM), produces such generalist time series models with minimal or no fine-tuning while exhibiting strong zero-shot performance. We evaluate TOTEM extensively over nearly 500 experiments on three commonly-studied time series tasks with real-world data: imputation (17 baselines, 12 datasets), anomaly detection (19 baselines, 25 datasets), and forecasting (14 baselines, 12 datasets). We conclude that TOTEM matches or outperforms existing state-of-the-art models in both the canonical specialist setting (i.e., training one model on one domain) as well as the generalist setting (i.e., training a single model on many domains), which demonstrates the efficacy of tokenization for general time series analysis. The open-source implementation is available here: https://github.com/SaberaTalukder/TOTEM; a video summary is available here: https://www.youtube.com/watch?v=OqrCpdb6MJk.
@@ 1 Introduction
Time series are a fundamental data modality, generalizing large classes of time-varying data from many domains, like weather phenomena, electrical grid activity, or traffic flow. Most commonly, time series analysis is first restricted to one such domain, then to a specific task, like imputation (Luo et al., 2018; 2019; Talukder et al., 2022), anomaly detection (Xu et al., 2021; He & Zhao, 2019), or forecasting (Wu et al., 2021; Woo et al., 2022), among others. Though these domains and tasks are quite distinct, a natural question is whether it is possible to design domain-agnostic models adaptable to any task. This question is the subject of our work.
@@
Generalist models are those trained on many data domains simultaneously (e.g., weather, electricity, traffic, etc.), while specialist models are those trained on a single time series domain (e.g., weather only), as shown in Figure 1A (Zhou et al., 2023; Wu et al., 2022; Nie et al., 2022). Both generalist and specialist models can be tested in two ways: in-domain testing, where a model is tested on the same domain(s) on which it was trained, and zero-shot testing, where it is tested on different domain(s) (see Figure 1B). Performing zero-shot testing on specialist models is not uncommon. For example, some works have studied zero-shot forecasting, where a forecaster trains on one dataset then predicts on a separate dataset (Zhou et al., 2023), or trains on a subset of channels (which we call sensors) from one dataset then forecasts zero-shot on the remaining sensors in the same dataset (Liu et al., 2023). However, we emphasize that both of the preceding examples are specialists, as they were trained on only one (or a subset of one) dataset. In contrast, our goal in this paper is instead the design of generalist models, which we evaluate in both the in-domain and zero-shot testing regimes.
Not only are most modern time series models specialists, they typically operate over patches (Zhou et al., 2023; Wu et al., 2022; Liu et al., 2023; Zhang & Yan, 2022; Nie et al., 2022; Li et al., 2019; Zhou et al., 2021; Wu et al., 2021) and are trained for only a single task (Das et al., 2023b; Ansari et al., 2024; Liu et al., 2023; Zhang & Yan, 2022; Zhou et al., 2021; Wu et al., 2021). Our core hypothesis is that many of the design decisions in prior works hinder the development of generalist models, and that by adopting practices more commonly used in language (Gage, 1994; Radford et al., 2018) and vision modeling (Van Den Oord et al., 2017; Esser et al., 2021; Rombach et al., 2022) we can boost the generalization performance of resulting time series models. While there exist works that train in an unsupervised manner (Yue et al., 2022; Yang & Hong, 2022; Tonekaboni et al., 2021; Barnum et al., 2020; Franceschi et al., 2019) or use discrete representations (Rabanser et al., 2020b; Van Den Oord et al., 2017; Lin et al., 2007), few works have explored the combination of generalist models and discrete representations over many tasks in a systematic manner (i.e., in both the in-domain and zero-shot testing regimes). Thus, the contributions of our work are twofold.
TOTEM. We develop Tokenized Time Series Embeddings (TOTEM), a simple tokenization method for time series data that employs a self-supervised pre-training stage to learn a fixed number of discrete tokens over a multi-domain corpus (Section 3.2). Surprisingly, we demonstrate that TOTEM is effective for solving a variety of downstream tasks in a domain-agnostic manner even though the tokens only encode the shape of univariate waveforms. This allows TOTEM to generically tokenize multivariate data of differing size by simply stacking collections of univariate tokens.
Comprehensive Experiments. We test our hypothesis extensively on three distinct tasks, each with their own datasets and baselines: imputation (17 baselines and 12 datasets), anomaly detection (19 baselines and 25 datasets), and forecasting (14 baselines and 12 datasets). We find that in the specialist settings, TOTEM matches or outperforms the performance of most state-of-the-art (SOTA) task-specific models, despite minimal or no task-specific design. Similarly, TOTEM also matches or outperforms SOTA generalist models. We conduct thorough ablations showing that discrete tokens outperform patches and that generalist training improves model performance independent of TOTEM’s modeling choices. Our experiments are some of the most extensive in the literature, comprising hundreds of seeded runs (see Sections 5 and 6).
@@ 2 Related Work
We categorize related works in three ways: whether they (i) study specialists or generalists, (ii) use patched or discrete data representations, and (iii) train and evaluate models for multiple distinct time series tasks. Unlike TOTEM, no prior works study the use of discrete data representations for training generalists across multiple tasks (see Table 1 for a comparison).
Specialist vs. Generalist Training. Historically, the specialist (i.e., single-domain) training paradigm is most common amongst time series models (Zhou et al., 2023; Wu et al., 2022; Nie et al., 2022; Zhang & Yan, 2022). These specialist models are primarily evaluated via in-domain testing, where the test set is a held-out set from the same training domain. Some concurrent and subsequent works have begun exploring generalist time series foundation models, including forecasters from Google and Amazon (Das et al., 2023b; Ansari et al., 2024). We compare to the concurrent MOMENT model (Goswami et al., 2024) in limited evaluations (see Tables 13 and 20) as it also studies multiple tasks, and find that TOTEM generally outperforms it.
Patched vs. Discrete Data Representations. In order to pass time series data to a downstream model, it is necessary to choose some latent data representation. As in ViTs (Dosovitskiy et al., 2020), the prevailing strategy is to patch time series data, either temporally (Liu et al., 2023; Zhang & Yan, 2022; Nie et al., 2022) or spatially Li et al. (2019); Zhou et al. (2021); Wu et al. (2021), then to linearly project the patches to some latent embedding on which a model like a transformer or MLP can operate. We emphasize that patched representations are dynamic in the sense that the embedding associated with the patch is determined entirely by the layers in the downstream model which project the patches to the embedding space. Therefore, patched representations are trained end to end.
Patching is fundamentally at odds with tokenization, wherein a fixed “vocabulary” of embeddings is determined before training the downstream model, which then operates on the fixed, tokenized representations. Tokenization (learned or otherwise) has been leveraged for training models in fields like language and vision modeling (Gage, 1994; Radford et al., 2018; Van Den Oord et al., 2017; Esser et al., 2021; Rombach et al., 2022). Some prior work in time series modeling has explored discrete representations using binning (Rabanser et al., 2020b; a; Lin et al., 2007) or quantization (Baevski et al., 2020; Van Den Oord et al., 2017; Oord et al., 2016) in domain- or task-specific ways. Inspired by the success of vector quantized variational autoencoders (VQVAEs) in both audio and vision (Van Den Oord et al., 2017; Esser et al., 2021; Rombach et al., 2022), we build on these works by showing that the VQVAE is also effective for learning discrete representations for general time series modeling.
Generalist Training Discrete Tokenization Multiple Tasks
Prior GPT2 (Zhou et al., 2023) x x
TiNet (Wu et al., 2022) x x
W2V2.0 (Baevski et al., 2020) x x
SAX (Lin et al., 2007) x
C/S TimesFM (Das et al., 2023b) x x
Chronos (Ansari et al., 2024) x
MOMENT (Goswami et al., 2024) x
TOTEM (Ours)
Table 1: Related Work Overview. TOTEM is designed for generalist training using discrete tokenization for any task. No prior and concurrent/subsequent (C/S) works study all three at once.
Time Series Tasks. Prior works on time series modeling study a variety of tasks, like forecasting, anomaly detection, imputation, and classification. Many prior and concurrent works focus on a single task (Zhang & Yan, 2022; Nie et al., 2022; Xu et al., 2021; Ansari et al., 2024; Das et al., 2023b), with a few exploring multiple specialist trained models on many tasks (Zhou et al., 2023; Wu et al., 2022). TOTEM is most closely related to concurrent works like MOMENT (Goswami et al., 2024), which are focused on generalist models which are effective on any one of the above tasks. For detail on each task, see Sections 3 and 4.